# ReUseIt: Synthesizing Reusable AI Agent Workflows for Web Automation

### v1

@@ 3.1. Evaluation: Reliability of web agents to execute common web interaction tasks
Figure 4 summarizes the experiment results. Across tasks, the $SR$s are: Task-Only ($27.9\%\pm 6.3\%$), Task + Success-Traces ($61.4\%\pm 16.1\%$), and Task + Magentic-UI Plan ($68.1\%\pm 15.1\%$). These results show that (1) although each task succeeded at least once, the agent’s performance is inconsistent across runs and the average $SR$ remains low; and (2) while successful execution traces can greatly improve the $SR$ for repeating the same or similar tasks, a gap remains before users can fully hand off web interaction tasks to agents.
Figure 4. Evaluation of web agents on six web automation tasks and their variations (avg $\pm$ std success rate).
3.2. Formative Study: Challenges for users to apply web agents for web automation
@@
  • Can users identify agent issues, how do they find issues, and what makes diagnosis challenging?
@@
Table 1. Participants’ self-reported frequency of LLM use, expertise in basic, intermediate, and advanced prompting (1=strongly disagree, 5=strongly agree), and time taken (secs) to find agent issues and provide guidance.
@@
[…]
During the study, after a brief introduction and example walkthrough ($\sim$10 mins), each participant was asked to inspect three web automation tasks that web agents failed to complete (we shuffled the task order and ensured that each task was examined by three different participants; $\sim$30 mins) and provide feedback to the agent to help it recover. In each task, participants were provided with slides containing the task description, the screen recording of agent actions, and the workflow the agent followed in the form of a list text block. The workflow was generated with Magentic-UI’s approach (Mozannar et al., 2025d). Participants were invited to verbally analyze why and how the agent failed to execute the tasks and edit the workflows on the slides. Finally, we conducted a semi-structured interview with each participant to get a deeper understanding of their user experience ($\sim$20 mins). More details on the formative study design can be found in Appendix A.2.
We performed thematic analysis (Braun and Clarke, 2006, 2012) using the qualitative data from interviews to code participants’ experience in spotting and fixing agent errors. We then performed a follow-up experiment to measure the web agent’s reliability in completing the tasks guided by these user-fixed workflows.
@@
User detection of agent issues. From the study, we found that the participants used different ways to look for agent issues. All participants were able to find agent issues by examining the screen recordings. The time taken by them to find issues is reported in Table 1. They frequently paused and replayed the recordings to check agent actions, such as how the agent handled pop-ups, applied filters, filled forms, and navigated pages. They described that the recordings helped them pinpoint problems (S1-P3) and understand the agent’s intent (S1-P6). Additionally, participants also cross-referenced the agent’s workflows (S1-P2: “I referenced the workflow to see the step breakdown and then came back to the video to check the actual actions.”), and used the workflows to infer potential causes of issues, such as missing instructions about applying filters in the workflow (S1-P4: “It [the agent] might not know the possible actions it can take… so it’s relying very much on keyword-based search.”).
The agent issues found by our participants were mostly low-level but critical mistakes, such as: (1) Filters not applied. The agent proceeded without applying all the requested filters. (2) Navigation challenges. The agent was stuck by pop-up banners, new pop-up UI elements like drop-down menus triggered by clicking a text field. (3) Random clicks/scroll. The agent clicked on something random or did not have a good awareness of the page layout, scroll position, or visible elements. (4) Ambiguity about intent. The agent was unclear about certain user instructions, such as “first/top/most recent”. (5) Partial actions. The agent forgot to finish clicking the search button or set all required parameters.
@@
User guidance to the agent. We observed that the participants used a combination of different strategies to provide guidance when the agent made mistakes, which can be summarized into five categories: (1) Improve the specificity of instructions. A common strategy was to add details to the workflow and specify each expected agent action (S1-P1, S1-P2, S1-P5). (2) Decompose instruction into clear steps. To ensure reliable execution, participants translated complex goals into explicit and clear steps for the agent to follow (S1-P2, S1-P6). (3) Add action verification for self-reflection. Participants highlighted building guardrails that confirm each critical action and branch when something goes wrong, so mistakes do not silently cascade across steps (S1-P1, S1-P3). (4) Provide action candidates for agent reference. Participants focused on reducing ambiguity by surfacing explicit action options the agent can choose from, such as available UI operations and their effects (S1-P4, S1-P6). (5) Include tips to handle webpage changes or common issues. To keep agent workflows resilient against varying webpage changes or recurring agent issues, participants embedded generic troubleshooting cues in the workflow (S1-P1, S1-P4).
Most participants focused on adding instruction details and clarifying required steps to reduce ambiguity. However, they expressed uncertainty regarding what level of specificity in the instruction was good for the agent to process and follow effectively (S1-P1, S1-P5). In addition, they mentioned that improving the specificity and clarity of instructions can be time-consuming, so they may perform the tasks manually instead (S1-P2: “I’d hope the AI could generate a prompt like this, and then it can do the task. Cuz I don’t really want to spend a lot of time on writing up this prompt since I would have already been able to do that myself.”). The time taken to provide guidance is summarized in Table 1.
@@
To understand the reliability of user-provided guidance (e.g., Figure 3 C), we conducted a follow-up experiment with the user-fixed workflows (Task + Magentic-UI Plan User), and the results are shown in Figure 4. The success rate in this figure was averaged across users, and Appendix A.3 reports the full results for each user. Across tasks and user-fixed workflows, Magentic-UI Plan User has improved the $SR$ by around $5.4\%$ compared to Magentic-UI Plan. From these results, the strategies that most aligned with higher and more consistent success were decomposing instructions into clear steps, and providing action candidates and recovery tips (S1-P4 and S1-P6 using these strategies achieved the top overall $SR$). A mix of adding specificity, action verification, and recovery tips by S1-P1 also yielded comparatively high $SR$s.
3.3. Reflection: Gaps in applying web agents for repetitive tasks
Our evaluation and formative study results show that while users can interactively work with web agents to complete tasks, their efforts spent on helping web agents might outweigh the benefit from agent automation. Specifically, we noticed the following three gaps for applying web agents for repetitive tasks.
Substantial time needed to understand issues. Across sessions, all participants devoted considerable time to reviewing screen recordings and monitoring the agent’s step-by-step actions to locate failures and reason about their causes. When actions were ambiguous, participants expressed uncertainty and frequently sought confirmation of their interpretations (S1-P1, S1-P2, S1-P3, S1-P5). This indicates that even careful inspection sometimes left them unsure whether their understanding of the issues was correct.
Significant effort to fix low-level mistakes. Examining the edits made to the workflows, we observed that participants consistently targeted the agent’s low-level action issues. Despite these efforts, user guidance improved the success rate by only a small margin according to our follow-up experiments. Participants also questioned whether they would want to invest this level of effort in their own tasks (S1-P2, S1-P5).
Users wanted their guidance to be effectively reused. Participants preferred web agents that they could reuse without having to re-specify guidance each time. For example, they pointed to tasks like flight booking (S1-P2), course registration (S1-P2), job search (S1-P3), social media/news/paper summarization (S1-P1, S1-P5, S1-P6), housing search (S1-P3, S1-P6), and price search (S1-P4, S1-P5). In these use cases, they were willing to invest in workflow improvements so long as those edits would carry forward to future runs and indeed help to improve the agent reliability.
@@ 4. System Design
  • DG1: Automatically mitigating low-level but critical issues. Low-level but critical agent issues (as introduced in Section 3.2) are often out of users’ expectations and can be hard to spot. Instead of relying on users to fix these errors hands-on, the web agent should avoid these errors and recover autonomously at runtime. Our approach is to let web agents automatically synthesize workflows from their successful and failed experiences to avoid repeating mistakes. We further develop an agent self-reflection mechanism for the agent to analyze web states to detect and recover from issues during execution (Section 4.1).
  • DG2: Surfacing issues to help users identify where and why agents fail. In cases where agents fail to find or fix minor issues, the system should provide users with additional information to understand where and why the failure occurs. Thus, users can act accordingly to take control and fix issues without replaying actions frame-by-frame to speculate the causes. As users provide feedback to help agents recover from issues, their feedback should be incorporated as part of the workflow to prevent error recurrence. We continuously update the workflows with user guidance and propagate them to future task executions (Section 4.2).
@@ 4.1. Automatic Workflow Synthesis (DG1)
To allow the agent to locate and correct issues on its own, we propose an automatic workflow synthesis approach to guide task execution. Inspired by the user strategies from the formative study, we synthesize workflows that provide detailed steps, webpage state checks, and recovery tips for the agent to follow the specific instructions, self-reflect to detect issues, and self-recover when issues occur.
In our approach, we obtain insights from both failed and successful agent executions. Failed executions can help expose common agent challenges that are then converted into pre- and post-condition checks around each action. For example, before navigating to a new page, the agent should verify that all required form fields are complete and correct; after entering a user-specified value in a text field, the field should display exactly that value. Successful executions, in turn, yield fallback actions, recovery strategies for steps where failures commonly occur, and workflow structure, the sequence of major steps, to which the derived condition checks and fallback actions are attached. The action details of a step, and its condition checks and fallback actions form a unit, which is concatenated with the rest of the steps/units to obtain the synthesized ReUseIt Workflow. Figure 5 illustrates this process.
@@
Multiple task executions. To synthesize the workflow, the first step is letting the web agent execute the tasks multiple times. Here, the tasks include the original task (directly from the user’s task description), as well as the variation tasks generated from the original task. The goal of adding variation tasks is to enable the synthesized workflow to generalize across closely related tasks. We adopt a systematic way to generate these variation tasks based on attribute, category, and website variations. These variations exercise different levels of agent generalizability: attribute variation requires value substitution on the same webpage; category variation involves tab/category switches within a webpage; website variation includes UI layout change on different websites. For example, for the flight search task “search the ¡flight type¿ ¡cabin type¿ ticket from ¡departure city¿ to ¡destination city¿ on ¡date¿ for ¡number of passengers¿ on ¡website¿”, attribute variations include ¡cabin type¿ (e.g., economy, business), ¡departure city¿, ¡destination city¿, ¡date¿, and ¡number of passengers¿; category variations include ¡flight type¿ (e.g., one-way, round-trip); website variations varies different flight search ¡website¿ (e.g., Booking.com, Expedia.com). To generate the variation tasks, ReUseIt uses the variation task generation prompt to guide the LLM to automatically instantiate task variants based on the original task. With the original task, the LLM identifies the values that can be substituted and groups these values into the attribute, category, and website variation categories. Within each category, the LLM needs to generate one variation task. We set the number of task executions to five for both the original and variation tasks, as this can lead to at least one successful execution for workflow synthesis based on our experiments.
@@
Execution guard synthesis. Multiple task executions yield failed and successful executions. For the failed attempts, we obtain the agent messages, which describe each agent action (e.g., navigated to a webpage, clicked a button, entered a value) and the resulting webpage state after each action. For successful attempts, we likewise capture the agent’s messages and additionally use Magentic-UI’s plan learning module (Mozannar et al., 2025a) to synthesize a high-level plan consisting of the major steps. This plan serves as the workflow structure of the ReUseIt Workflow. The prompt for plan learning was borrowed from Magentic-UI without modification. Leveraging the messages and plans from multiple task executions, we synthesize the execution guards, condition checks and fallback actions, to augment the workflow structure. First, ReUseIt uses the condition check synthesis prompt to guide the LLM to extract agent challenges from failed execution messages and convert these challenges into conditions that should be met. From error messages, such as “failed to, didn’t, or couldn’t” perform a certain action because “a button is inactive, a text field can’t be located, or a page doesn’t load,” the LLM extracts the error pattern and convert it into a pre- or post-condition check for this action in the form of “Before/After doing ¡action¿, ensure ¡condition¿ is met.” Next, ReUseIt assembles the fallback action synthesis prompt to ask the LLM to find agent actions from successful execution messages at failed steps and turn these successful experiences into recovery strategies. Based on the description of agent actions, such as “I clicked, I navigated to, or I typed ¡UI element¿,” the LLM turns these actions into retry strategies, such as “Retry performing ¡action¿ by clicking, navigating to, or typing ¡UI element¿.” Finally, ReUseIt assembles the workflow synthesis prompt to insert the synthesized condition checks and fallback actions into the corresponding step in the workflow structure to form the ReUseIt Workflow.
@@
Remarks. Our current approach allows a parallel workflow synthesis from both successful and failed experiences simultaneously. This method can be altered into a sequential alternative if the successful execution can be quickly obtained, e.g., within 3 runs. Based on this successful execution, we can obtain the workflow structure and iterate to add condition checks and fallback actions with further executions. For tasks that are challenging for the agent to easily obtain a successful execution, the parallel approach is more efficient as users do not need to wait long for a successful run. We used the parallel approach in our implementation to accommodate tasks with multiple difficulty levels.

### v2

@@ 3.1. Preliminary Evaluation: Reliability of web agents to execute common web automation tasks
Figure 4 summarizes the experiment results. Across tasks, the $SR$s are: Task-Only ($27.9\%\pm 6.3\%$), Task + Success-Traces ($61.4\%\pm 16.1\%$), and Task + Magentic-UI Plan ($68.1\%\pm 15.1\%$). In addition, we report the mean $SR$s with 95% confidence intervals of each task and across all tasks in Appendix A.2. These results show that (1) although each task succeeded at least once, the agent’s performance is inconsistent across runs and the average $SR$ remains low; and (2) while successful execution traces and plans can greatly improve the $SR$ for repeating the same or similar tasks, a gap remains before users can fully hand off web interaction tasks to agents.
Figure 4. Evaluation of web agents’ success rate on six web automation tasks and their variations (avg $\pm$ std success rate).
This figure shows six bar charts with task success rates for the web automation tasks.
3.2. Formative Study: Challenges for users to apply web agents in web automation tasks
@@
  • Can users identify the issues of web agents, how do they find issues, and what makes diagnosis challenging?
@@
Table 1. Participants’ self-reported frequency of LLM use, expertise in basic, intermediate, and advanced prompting (1=strongly disagree, 5=strongly agree), and time taken to find agent issues and provide guidance.
@@
S1-P1 Daily 5 5 5 $923$ s $988$ s
S1-P2 Daily 5 4 4 $952$ s $997$ s
S1-P3 Daily 5 5 4 $898$ s $858$ s
S1-P4 Daily 5 5 5 $867$ s $902$ s
S1-P5 Daily 4 4 2 $973$ s $933$ s
S1-P6 Daily 5 5 5 $885$ s $932$ s
This table summarizes the demographic information of the formative study participants, as well as the time they took to find agent issues and offer guidance to help the agent recover from issues during the study.
During the study, after a brief introduction and example task walkthrough ($\sim$10 mins), each participant was asked to inspect three web automation tasks that web agents failed to complete (we shuffled the task order and ensured that each task was examined by three different participants; $\sim$30 mins) and offer guidance to the agent to help it recover. In each task, participants were provided with slides containing the task description, the screen recording of agent actions, and the workflow the agent followed in the form of a list text block. The workflow was generated with Magentic-UI’s approach (Mozannar et al., 2025a). Participants were invited to verbally analyze why and how the agent failed to execute the tasks and edit the workflows on the slides. Finally, we conducted a semi-structured interview with each participant to get a deeper understanding of their user experience ($\sim$20 mins). More details of the formative study materials can be found in Appendix B.1.
We performed thematic analysis (Braun and Clarke, 2006, 2012) using the qualitative data from interviews to code participants’ experience in spotting and fixing agent errors. We then performed a follow-up experiment to evaluate the web agent’s performance in completing the tasks guided by these user-fixed workflows.
@@
User detection of agent issues. From the study, we found that the participants used different ways to identify agent issues. All participants were able to find agent issues by examining the screen recordings. The time taken by them to find issues is reported in Table 1. They frequently paused and replayed the recordings to check agent actions, such as how the agent handled pop-ups, applied filters, filled forms, and navigated pages. They described that the recordings helped them pinpoint problems (S1-P3) and understand the agent’s intent (S1-P6). Additionally, participants also cross-referenced the agent’s workflows (S1-P2: “I referenced the workflow to see the step breakdown and then came back to the video to check the actual actions.”), and used the workflows to infer potential causes of issues, such as missing instructions on how to apply filters (S1-P4: “It [the agent] might not know the possible actions it can take… so it’s relying very much on keyword-based search.”).
The agent issues found by our participants were mostly low-level but critical mistakes, such as: (1) Filters not applied. The agent proceeded without applying all the requested filters. (2) Navigation challenges. The agent was stuck by pop-up banners, new pop-up UI elements like drop-down menus triggered by clicking a text field. (3) Random clicks/scroll. The agent clicked on something random or did not have a good awareness of the page layout, scroll position, or visible elements. (4) Ambiguity about intent. The agent was unclear about certain user instructions, such as “first/top/most recent”. (5) Partial actions. The agent failed to finish clicking the search button or set all required parameters.
@@
User guidance to the agent. When the agent made mistakes, we observed that the participants used a combination of different strategies to provide guidance, which can be summarized into five categories: (1) Improve the specificity of instructions. A common strategy was to add details to the workflow and specify each expected agent action (S1-P1, S1-P2, S1-P5). (2) Decompose instructions into clear steps. To ensure reliable execution, participants translated complex instructions into explicit and clear steps for the agent to follow (S1-P2, S1-P6). (3) Add action verification for self-reflection. Participants highlighted building guardrails that confirm each critical action before proceeding, and branching to take alternative actions when something goes wrong so mistakes do not silently cascade (S1-P1, S1-P3). (4) Provide action candidates for agent reference. Participants focused on reducing ambiguity by surfacing explicit action options the agent can choose from, such as available UI operations and their effects (S1-P4, S1-P6). (5) Include tips to handle webpage changes or common issues. To keep agent workflows resilient against varying webpage changes or recurring agent issues, participants embedded generic troubleshooting cues in the workflow (S1-P1, S1-P4).
Most participants focused on adding instruction details and clarifying required steps to reduce ambiguity. However, they expressed uncertainty regarding what level of specificity in the instruction was good for the agent to effectively process and follow (S1-P1, S1-P5). In addition, they mentioned that improving the specificity and clarity of instructions can be time-consuming, so they may perform the tasks manually instead (S1-P2: “I’d hope the AI could generate a prompt like this, and then it can do the task. Cuz I don’t really want to spend a lot of time on writing up this prompt since I would have already been able to do that myself.”). The time taken for participants to provide guidance is summarized in Table 1.
@@
To understand the effectiveness of user-provided guidance (e.g., Figure 3 C), we conducted a follow-up experiment with the user-fixed workflows, and the results are shown in Figure 4 (Task + Magentic-UI Plan User). The success rate in this figure was averaged across users, and Appendix B.2 reports the full results for each user. Across tasks and user-fixed workflows, Magentic-UI Plan User has improved the $SR$ by around $5.4\%$ compared to Magentic-UI Plan. From these results, the strategies that most aligned with higher and more consistent success were decomposing instructions into clear steps, and providing action candidates and recovery tips (S1-P4 and S1-P6 using these strategies achieved the top overall $SR$). A mix of adding specificity, action verification, and recovery tips by S1-P1 also yielded comparatively high $SR$s.
3.3. Reflection: Gaps in applying web agents for repetitive web tasks
Our preliminary evaluation and formative study results show that while users can interactively work with web agents to complete tasks, their efforts spent on helping web agents might outweigh the benefits from agent automation. Specifically, we noticed the following three gaps for users to apply web agents in repetitive tasks.
Substantial time needed to understand issues. All our participants devoted considerable time to reviewing screen recordings and monitoring the agent’s step-by-step actions to locate failures and reason about their causes. When actions were ambiguous, participants expressed uncertainty and frequently sought confirmation of their interpretations (S1-P1, S1-P2, S1-P3, S1-P5). This indicates that even careful inspection sometimes left them unsure whether their understanding of the issues was correct.
Significant effort to fix low-level but critical mistakes. Examining the edits made to the workflows, we observed that participants consistently targeted the agent’s low-level but critical action issues. Despite these efforts, user guidance improved the success rate by only a small margin according to our follow-up experiments. Participants also questioned whether they would want to invest this level of effort in their own tasks (S1-P2, S1-P5).
Users wanted their guidance to be effectively reused. Participants preferred web agents that they could reuse without having to re-specify guidance each time. For example, they pointed to tasks like flight booking (S1-P2), course registration (S1-P2), job search (S1-P3), social media/news/paper summarization (S1-P1, S1-P5, S1-P6), housing search (S1-P3, S1-P6), and product price search (S1-P4, S1-P5). In these use cases, they were willing to invest in workflow improvements so long as those edits would carry forward to future runs and indeed help to improve the agent reliability.
@@ 4. System Design
  • DG1: Automatically mitigating low-level but critical agent issues. Low-level but critical agent issues (as introduced in Section 3.2) are often out of users’ expectations and can be hard to spot. Instead of relying on users to fix these errors hands-on, the web agent should avoid these errors and recover autonomously at runtime. Our approach is to let web agents automatically synthesize workflows from their successful and failed experiences to avoid repeating mistakes. We further develop an agent self-reflection mechanism for the agent to analyze web states to detect and recover from issues during execution (Section 4.1).
  • DG2: Surfacing issues to help users identify where and why agents fail. In cases where agents fail to find or fix minor issues, the system should provide users with additional information explaining where and why the failure occurs. Thus, users can act accordingly to take control and fix issues without replaying agent actions frame-by-frame to speculate the causes. As users provide feedback to help agents recover from issues, their feedback should be incorporated as part of the workflow to prevent error recurrence. We continuously update the workflows with user guidance and propagate them to future task executions (Section 4.2).
@@ 4.1. Automatic Workflow Synthesis (DG1)
To allow the web agent to locate and correct issues on its own, we propose an automatic workflow synthesis approach to guide task execution. Inspired by the user strategies from the formative study, we synthesize workflows that provide detailed steps, webpage state checks, and recovery tips for the agent to follow specific instructions, self-reflect to detect issues, and self-recover when issues occur.
In our approach, we synthesize workflows from both failed and successful agent executions. Failed executions can help expose common agent challenges that are then converted into pre- and post-condition checks around each action. For example, before navigating to a new page, the agent should verify that all required form fields are complete and correct; after entering a user-specified value in a text field, the field should display exactly that value. Successful executions, in turn, yield fallback actions, the recovery strategies for steps where failures commonly occur, and workflow structure, the sequence of major steps to which the derived condition checks and fallback actions are attached. The action details of a step and the corresponding condition checks and fallback actions form a unit, which is concatenated with the rest of the steps/units to obtain the synthesized ReUseIt Workflow. Figure 5 illustrates this workflow synthesis process.
@@
This figure shows a flowchart for the automatic workflow synthesis pipeline. Starting from the left, under ”Task Variations”, there are four task variations. These variations are executed multiple times to obtain failed and successful attempts under ”Multiple Task Executions”. These attempts are used to synthesize execution guards, as shown under ”Execution Guard Synthesis”. Specifically, failed attempts reveal challenges faced by the agent and are turned into condition checks; successful attempts identify successful experience and are converted into fallback actions. Lastly, these synthesized execution guards are concatenated with the workflow structure to form the synthesized ReUseIt workflow.
@@
Multiple task executions. To synthesize the workflow, the first step is letting the web agent execute web tasks multiple times. Here, the tasks include the original task (directly from the user’s task description), as well as the variation tasks generated from the original task. The goal of adding variation tasks is to enable the synthesized workflow to generalize across closely related tasks. We adopt a systematic way to generate these variation tasks based on attribute, category, and website variations. These variations exercise different levels of agent generalizability: attribute variation requires value substitution on the same webpage; category variation involves tab/category switches within a webpage; website variation includes UI layout change on different websites. For example, for the flight search task “search the ¡flight type¿ ¡cabin type¿ ticket from ¡departure city¿ to ¡destination city¿ on ¡date¿ for ¡number of passengers¿ on ¡website¿”, attribute variations include ¡cabin type¿ (e.g., economy, business), ¡departure city¿, ¡destination city¿, ¡date¿, and ¡number of passengers¿; category variations include ¡flight type¿ (e.g., one-way, round-trip); website variations varies different flight search ¡website¿ (e.g., Booking.com, Expedia.com). To generate the variation tasks, ReUseIt uses the variation task generation prompt (Appendix C.1) to guide the LLM to automatically instantiate task variants. With the original task, the LLM identifies the values that can be substituted and groups these values into the attribute, category, and website variation categories. Within each category, the LLM needs to generate one variation task. We set the number of task executions as five for both the original and variation tasks, as this can lead to at least one successful execution for workflow synthesis based on our experiments.
@@
Execution guard synthesis. Multiple task executions yield failed and successful executions. For the failed attempts, we obtain the agent messages, which describe each agent action (e.g., navigated to a webpage, clicked a button, entered a value). For successful attempts, we likewise capture the agent’s messages and additionally use Magentic-UI’s plan learning module (Mozannar et al., 2025b) to synthesize a high-level plan consisting of the major steps. This plan serves as the workflow structure of the ReUseIt Workflow. The prompt for plan learning was borrowed from Magentic-UI  [^1] https://github.com/microsoft/magentic-ui/blob/main/src/magentic_ui/learning/learner.py. Leveraging the messages and plans from multiple task executions, we synthesize the execution guards—condition checks and fallback actions—to augment the workflow structure. First, ReUseIt uses the condition check synthesis prompt (Appendix C.2) to guide the LLM to extract agent challenges from failed execution messages and convert these challenges into conditions that should be met. From error messages, such as “failed to, didn’t, or couldn’t” perform a certain action because “a button is inactive, a text field can’t be located, or a page doesn’t load,” the LLM extracts the error pattern and convert it into a pre- or post-condition check for this action in the form of “Before/After doing ¡action¿, ensure ¡condition¿ is met.” Next, ReUseIt assembles the fallback action synthesis prompt (Appendix C.3) to ask the LLM to find agent actions from successful execution messages at failed steps and turn these successful experiences into recovery strategies. Based on the description of agent actions, such as “I clicked, I navigated to, or I typed ¡UI element¿,” the LLM turns these actions into retry strategies, such as “Retry performing ¡action¿ by clicking, navigating to, or typing ¡UI element¿.” Finally, ReUseIt assembles the workflow synthesis prompt (Appendix C.4) to insert the synthesized condition checks and fallback actions into the corresponding step in the workflow structure to form the ReUseIt Workflow.
@@
Remarks. Our current approach allows a parallel workflow synthesis from both successful and failed experiences simultaneously. This method can be altered into a sequential alternative if the successful execution can be quickly obtained, e.g., within 3 runs. Based on this successful execution, we can obtain the workflow structure and iterate to add condition checks and fallback actions with further executions. For tasks that are challenging for the agent to easily obtain a successful execution, the parallel approach is more efficient since users do not need to wait long for a successful run. We used the parallel approach in our implementation to accommodate tasks with multiple difficulty levels.